theyadoreshayyy:

Quinn is flying a kite. The angle of elevation formed by the kite string and the ground is 46°, and the kite string forms a straight segment that is 80 feet long. Explain how to find the distance between the ground and the kite. Include a description of the triangle you drew to help you solve, including the variables and measurements you assigned to each side and angle. Round your answer to the nearest foot.

darkknight:

as a visual

darkknight:

if not clear, that is 80 feet and the angle is 46 degrees. Finding the distance b/w the ground and the height involves trig, which trig ratio can you use? 

theyadoreshayyy:

tangent ?

darkknight:

You are given the length of the hypotenuse, and need to solve for the side opposite the angle, opposite/hypotenuse is?

theyadoreshayyy:

sine

darkknight:

So sin(46) = x/80 what do we do next?

theyadoreshayyy:

x/80=sin(46 degrees)

darkknight:

hold up, we are isolating for x you didn't do anything but switch the left and right sides above, I meant how would we solve for x

darkknight wrote:
So sin(46) = x/80 what do we do next?

theyadoreshayyy:

80x/80 =80sin(46 degrees)

theyadoreshayyy:

am i right ? or was that wrong ?

darkknight:

@theyadoreshayyy wrote:
80x/80 =80sin(46 degrees)
which equals x=80sin(46) plug into calculator and you get your height, in all the triangle we used is the triangle I drew above, the steps we just showed, and we used the variable x as the height which is what we solved for

theyadoreshayyy:

how would i do that ?

darkknight:

how to plug into a calculator?

theyadoreshayyy:

yes

darkknight:

Use fingers, and click the respective buttons..? Its not that hard put in 80*sin(46) into a calculator

theyadoreshayyy:

i got 3,680.

theyadoreshayyy:

@darkknight

darkknight:

What are you plugging, the hypotenuse is the longest side, the height can't be greater than 80, how do you get 3 thousand?

theyadoreshayyy:

no im not on windows